OpenStudy (anonymous):

A ball on a string moves around a complete circle, once a second, on a frictionless, horizontal table. The tension in the string is measured to be 6.0N . What would the tension be if the ball went around in only half a second?

OpenStudy (anonymous):

visit http://hyperphysics.phy-astr.gsu.edu/hbase/waves/string.html \[f=\sqrt{T/2ml}\] making '2ml' the subject of the formula \[2ml=T/f ^{2}\] then we can say \[T1/f1 ^{2}= T2/f2^{2}\] Then you can go on from there... I think the answer would be \[1.5Newton\]

OpenStudy (anonymous):

vorex, are you looking at different resonant frequency? o.O cuz f is frequency.. i think.

OpenStudy (anonymous):